39 The Keynesian consumption function is C = 15 + 0:6Y ; combined with C + 10 = Y (I = 10), an increase in government spending causes Y to rise by a multiplier of

a. 1.5.

b. 2.0.

c. 2.5.

d. 0.

59 If income is the "interest áow" on Wealth, then

a. income is greater than Wealth;

b. income is equal Wealth;

c. income is less than Wealth;

d. income is unrelated to Wealth.
Answer #1

Ans.39- 2.5

Multiplier = 1/1- mpc = 1/1-0.6 = 1/0.4 =2.5

Hence correct answer us C.

For question no. 59 I am unable to understand the meaning of word written after interest in "interest áow" . Write its meaning in comment box so that I can answer question no.59 too.